OpenStudy (anonymous):

evaluate this integral integral (1-cos t/2)^2 sin t/2 dt u= 1-cos t/2

myininaya (myininaya):

if u=1-cos(t/2) then du=1/2 *sin(t/2) dt or 2du=sin(t/2) dt

myininaya (myininaya):

so now we have \[2\int\limits_{}^{}u^2du\]

OpenStudy (anonymous):

thats the part i don't get

OpenStudy (anonymous):

how to get du

myininaya (myininaya):

so you are having trouble with derivatives?

myininaya (myininaya):

the derivative of 1 is 0 and the derivative of -cos(t/2) is 1/2(-1)(-sin{t/2})=1/2 sin(t/2)

myininaya (myininaya):

u is a function of t so we have du/dt=1/2*sin(t/2) so by multipling dt on both sides we have du

OpenStudy (anonymous):

ok i understand now thanks!